Let f (x)= x²-4 for x<3



7 for x=3



2x+4 for x>3



a.)f(0)=



b.)f(3)=



c.)f(5)=



d )lim f(x) as x approaches to 0=



e.)lim f(x) as x approaches to 3 from left=



f.)lim f(x) as x approaches to 3 from right=




1
Expert's answer
2022-03-01T15:45:24-0500

a) "f(0)=0^2-4=-4"


b) "f(3)=7"


c) "f(5)=2*5+4=14"


d) "\lim_{x\to 0}=f(0)=0^2-4=-4"


e) "\lim_{x\to 3^-}=\lim_{x\to3^-}(3^2-4)=5"


f) "\lim_{x\to 3^+}=\lim_{x\to 3^+}(2*3+4)=10"